If a particle is thrown with velocity more than 20m/s vertically upward, then the distance travelled by the particle in last second of it`s ascent is... A) g (B) g/2. (C) g/4. (D) g/8

Grade:11

1 Answers

Samyak
6 years ago
Speed of projection is 20m/sThus time of flight is 2u/g. T=4 secDistance in 2 th second is= u+a(2n-1)/2Thus distance is 20-g(3)/220-15=5g/2